# Hardware-inspired Continuous Variables Quantum Optical Neural Networks

### v1

@@ 4.5 Generalization to ladder polynomials
This observation directly connects the introduced framework to perturbative implementations of weak non-Gaussian gates such as cubic or Kerr interactions lloyd1999quantum; braunstein2005quantum; yanagimoto2020engineering. A short-time evolution under a non-Gaussian Hamiltonian $\hat{H}_{\mathrm{ng}}\in\{\hat{H}_{\mathrm{cubic}},\hat{H}_{\mathrm{Kerr}},\dots\}$ can be written as a power series
@@
where each term is a polynomial in the bosonic ladder operators. In the weak-gate regime (fast-gate), retaining only the leading orders in $t$ precisely yields the kind of polynomial structure captured by Eq. (33). The ansatz proposed in this work therefore provides an exact, non-perturbative evaluation of the homodyne statistics associated with such perturbative non-Gaussian gates, and can in particular reproduce and systematize the continuous-variable neural network constructions proposed by Lloyd and co-workers lloydcvqnn and by Siopsis and collaborators siopsis2023exprealcvqnn within a single unified framework without truncation.
From a hardware perspective, this generality implies that this scheme is not restricted to photonic platforms employing heralded photon subtraction. Whenever a platform can implement weak non-Gaussian interactions whose short-time expansion gives rise to polynomials in ladder operators, the same theoretical machinery and universality results apply. Examples include levitated opto-mechanical systems bemani2025heralded; muffato2025generation; marti2024quantum, superconducting microwave circuits blais2021circuit; kounalakis2018tuneable; hu2023fast, and other bosonic architectures where cubic, Kerr, or more general non-Gaussian Hamiltonians are available. In this sense, the photon-subtraction implementation studied here should be viewed as a particularly mature and experimentally feasible instance of a broadly applicable continuous-variable quantum optical neural network paradigm.
@@ 5.1 Curve fitting
[b] 0.34 Refer to caption {subcaptionblock}[b] 0.335 Refer to caption {subcaptionblock}[b] 0.31 Refer to caption
Figure 3: Curve fitting training of NN-mode single-layer ($L=1$) QONNs with photon subtraction $\hat{a}$ in different modes trained over 100-sample noisy datasets of exponential, polynomial and oscillatory functions. Left: Gaussian, single-neuron and two-neuron QONNs results for $f(x)=x^{5}$ with $x\in(-3,3),\ \epsilon=5$. Center: one, two and three-neuron QONNs results for $f(x)=2\cosh(x)$ with $x\in(-5,5),\ \epsilon=3$. Right: one, two, three and four-neuron QONNs results for $f(x)=\sin(3x)+\cos(5x)$ with $x\in(-1,2.5),\ \epsilon=0.1$.
@@
  • Synthetic training dataset of a target function $f(x)$ constructed by randomly sampling it within a specific range of $x$ from a uniform distribution with some noise level $\epsilon$.
@@
Fig. 3 shows the different single-layer QONN trainings indicating the modes where a photon subtraction is present. The left plot proves that, when there are no subtractions—i.e. QONN with only Gaussian operators—, learning is limited to a linear fit due to the lack of activations. Overall, these results show that adding more neurons—arising from the amount of subtracted modes—in a single layer makes the fitting error tend to zero regardless the complexity of the function, in line with the UAT.
@@ 5.2.1 Moons and circles classification
To test the classification skills of the QONN, two toy-example tasks to be tackled are the 2D moons and circles classification, where the target is to learn a nonlinear boundary that tells apart two different classes of points arranged in a continuous two-dimensional grid.
@@
According to the previously defined features, $N=2$ modes are sufficient to perform the tasks since both problems are binary classifications and the input space is two-dimensional and real-valued.
In Fig. 4, upper row shows how a Gaussian QONN trains up to a linear decision boundary for the moons problem, while a single-layer QONN with one neuron is able to fully discern between the two classes with a curved decision boundary, achieving perfect accuracy. On the other hand, the circles problem needs a radial decision boundary in order to be solved. The lower row of Fig. 4 demonstrates that a $2$-mode QONN with just one neuron fails to capture such a boundary where a two-neuron QONN succeeds with perfect precision.
@@
Figure 4: Moons classification (upper row): two interleaved categories of points. Upper-left: Linear decision boundary with 86.8% accuracy of a Gaussian QONN (no subtractions). Upper-right: S-like decision boundary with 100% of accuracy of a single-layer single-subtraction QONN. Circles classification (lower row): two concentric categories of points. Lower-left: highly-curved decision boundary with 93% of accuracy of a single-layer single-subtraction QONN. Lower-right: radial decision boundary with 100% of accuracy of a single-layer two-subtractions QONN (same accuracy as a 2-layer QONN with one subtraction per layer).
@@ 5.2.2 MNIST classification
  • Real-valued inputs: Autoencoder compressing the $28\times 28$ images to a continuous latent space of $3$ real-domain dimensions.
@@
Figure 5: 5-category MNIST classification showing per-class accuracy (opaque lines) and average precision (transparent horizontal lines) for different QONN models with an input pre-processing of 3-dimensional latent space.
@@ 5.3 Non-Gaussian gate synthesis
This admits to train the proposed QONN to match the statistical moments of any complex non-Gaussian gate acting over a particular set of quantum states. This means that the trained QONN block would be able to replicate the action of the chosen non-Gaussian operator upon the target quantum states using physically available quantum optical components, similar to yukawa2013.
@@ 6 Discussion
An important conceptual difference between this QONN proposal and earlier continuous-variable neural network architectures lloydcvqnn; siopsis2023exprealcvqnn is the role of Fock-space truncation. In those works, all quantities of interest are ultimately computed in a truncated Fock basis. Formally, this renders the scheme equivalent to a—very large—discrete-variable model: the dynamics is projected onto a finite-dimensional subspace, classical simulation inherits the unfavorable scaling of the truncated Hilbert space, and the resulting figures of merit are not directly expressed in terms of experimentally accessible quadrature statistics. In contrast, the presented approach is based on a genuinely continuous-variable ansatz and an exact evaluation of homodyne moments, without any photon-number cutoff, by combining Bogoliubov transformations with generalized Wick–Isserlis theorems. This provides a closed-form description of the effective activation functions and a classical simulation framework whose cost scales with the number of modes, layers, and photon subtractions, rather than with an artificially imposed Fock cutoff. For the tasks studied here, substantial practical speedup is observed over state-of-the-art Fock-space simulators such as QuTiP and Strawberry Fields.
These differences have direct implications for both expressivity and experimental feasibility. For instance, in the architecture of siopsis2023exprealcvqnn, the network depth reaches up to twelve layers with three photon-subtraction events per layer. At realistic squeezing levels and detection efficiencies, such a configuration would lead to heralding rates that are exceedingly low—on the order of one successful detection per day—, implying training times of years even under optimistic assumptions. To keep the numerics tractable, the analysis therefore relies on a photon-number cutoff, typically $n_{\max}=6$, reintroducing the drawbacks of truncation at the simulation level. In the exact framework introduced, by contrast, a single layer with only a few photon-subtracted modes is already sufficient to match or surpass the performance reported siopsis2023exprealcvqnn on comparable tasks, while remaining compatible with realistic heralding probabilities on current photonic platforms. This highlights that a shallow, analytically tractable continuous-variable quantum optical neural network can simultaneously achieve high expressive power, efficient classical simulability, and genuine experimental viability without resorting to Fock space truncation.
@@ 7 Code availability
Since the QONN outputs are computed as a sum of traces and each of such traces can be individually obtained via Wick–Isserlis expansion of Eq. (LABEL:perf_match)—a term-independent sum of products—, high-performance environments are perfect candidates for the classical simulation of the model. Therefore, QuaNNTO library has been implemented using the large-scale high-performance JAX package jax2018github together with the Basinhopping method from SciPy library that embeds the L-BFGS-B optimization algorithm with default settings, allowing exhaustive simulations of gradient-based QONN trainings with distributed and parallel computation.
@@ Appendix A Nonlinearity of the creation operator in phase-space
Eq. (A2) is equivalent to Laguerre polynomials defined as creation operators over coherent states [agarwal1991nonclassical], whose expansion leads to the general expression
@@
$\langle\hat{x}(\alpha,r)\rangle=\sqrt{2}\Re\Big(\underbrace{e^{r}\,\alpha}_{\text{Linear map}\alpha}\;+\;\underbrace{e^{r}\,\alpha\,\frac{1+e^{r}\sinh r}{\alpha^{2}e^{2r}+\cosh^{2}r}}_{\text{Nonlinear bump}}\Big).$ (A6)
Figure 7: Nonlinearity induced by photon addition operator on Gaussian states in $\hat{x}$ quadrature. Right: $\langle\hat{x}(\alpha)\rangle_{n}$ from Eq. (A3) for $n=1..4$ photon additions with range of the real displacement $\alpha\in(-2,2)$. Left: $\langle\hat{x}(\alpha,r)\rangle$ from Eq. (A6) for different values of squeezing in position ($r\geq 0$) and momentum ($r\leq 0$).
From the explicit shapes of Eq. (A3) shown in left Fig. 7, one can clearly notice the smooth nonlinear sigmoid-like effect of the photon addition operations—bounded and non-polynomial—over a real coherent state when the displacement is relatively small. However, the effect tends to be linear when the displacement is more pronounced, making it a saturated nonlinearity.
The effect of the photon addition is similar when the initial coherent state is squeezed previously to the addition, as seen in Eq. (A6). The activation function is also non-polynomial and smooth. The right plot of Fig. 7 shows how squeezing momentum quadrature pronounces the nonlinearity induced by the photon creation in the position quadrature mean, while squeezing position tends to linearize such effect.
@@ Appendix B Exact classical simulation of the CV-QONN model
To classically simulate the QONN from Fig. 1 using the Wick–Isserlis expansion of Eq. (LABEL:perf_match), all the Gaussian operators from the layers have to be dragged to the beginning of the circuit for the state to remain Gaussian and, afterwards, apply all the non-Gaussianity made of creation and annihilation operators.
This is achieved by the Bogoliubov unitary action of the general Gaussian operator $\hat{G}$ of the form of Eq. (9) over the annihilation operator [ulysse2021classical]
@@
$\rho_{NG}=\frac{\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\ \rho_{G}\ \hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}}{Tr\big[\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\ \rho_{G}\ \hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}\big]},$ (B7)
@@
$\begin{split}\tilde{f}(\bm{\alpha})=Tr\left[\hat{x}_{1}\rho_{NG}\right]=\frac{Tr\left[\frac{(\hat{a}_{1}+\hat{a}^{\dagger}_{1})}{\sqrt{2}}\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\ \rho_{G}\ \hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}\right]}{Tr\big[\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\ \rho_{G}\ \hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}\big]}=\frac{\sqrt{2}\cdot\Re\Big(Tr\left[\hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}\hat{a}_{1}\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\rho_{G}\right]\Big)}{Tr\big[\hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\rho_{G}\big]},\end{split}$ (B9)
@@
Each of these trace expressions has the form of Eq. (LABEL:perf_match) and is computed by Wick–Isserlis using the set $\mathcal{P}$ of all loop perfect matchings of the ladder operators in the expression. For an expression with $M$ ladder operators, the total number of different matchings increases combinatorially, given by Eq. (11).
Note that additionally, this needs to be applied for the normalization factor of the QONN’s final quantum state $\rho_{NG}$ of Eq. (B7) with $O=1$ in Eq. (B10).

### v2

@@ 4.5 Generalization to ladder polynomials
This observation directly connects the introduced framework to perturbative implementations of weak non-Gaussian gates such as cubic or Kerr interactions Lloyd and Braunstein (1999); Braunstein and Van Loock (2005); Yanagimoto et al. (2020). A short-time evolution under a non-Gaussian Hamiltonian $\hat{H}_{\mathrm{ng}}\in\{\hat{H}_{\mathrm{cubic}},\hat{H}_{\mathrm{Kerr}},\dots\}$ can be written as a power series
@@
where each term is a polynomial in the bosonic ladder operators. In the weak-gate regime (fast-gate), retaining only the leading orders in $t$ precisely yields the kind of polynomial structure captured by Eq. (33). The ansatz proposed in this work therefore provides an exact, non-perturbative evaluation of the homodyne statistics associated with such perturbative non-Gaussian gates, and can in particular reproduce and systematize the continuous-variable neural network constructions proposed by Lloyd and co-workers Killoran et al. (2019a) and by Siopsis and collaborators Bangar et al. (2023) within a single unified framework without truncation.
From a hardware perspective, this generality implies that this scheme is not restricted to photonic platforms employing heralded photon subtraction. Whenever a platform can implement weak non-Gaussian interactions whose short-time expansion gives rise to polynomials in ladder operators, the same theoretical machinery and universality results apply. Examples include levitated opto-mechanical systems Bemani et al. (2025); Muffato et al. (2025); Marti et al. (2024), superconducting microwave circuits Blais et al. (2021); Kounalakis et al. (2018); Hu et al. (2023), and other bosonic architectures where cubic, Kerr, or more general non-Gaussian Hamiltonians are available. In this sense, the photon-subtraction implementation studied here should be viewed as a particularly mature and experimentally feasible instance of a broadly applicable continuous-variable quantum optical neural network paradigm.
@@ 5.1 Curve fitting
[b] 0.33 Refer to caption {subcaptionblock}[b] 0.33 Refer to caption {subcaptionblock}[b] 0.33 Refer to caption
Figure 3: Curve fitting training of NN-mode single-layer ($L=1$) QONNs with photon subtraction $\hat{a}$ in different modes trained over 100-sample noisy datasets of exponential, polynomial and oscillatory functions. Left: Gaussian, single-neuron and two-neuron QONNs results for $f(x)=x^{5}$ with $x\in(-3,3),\ \epsilon=5$. Center: one, two and three-neuron QONNs results for $f(x)=2\cosh(x)$ with $x\in(-5,5),\ \epsilon=3$. Right: two, three and four-neuron QONNs results for $f(x)=\sin(3x)+\cos(5x)$ with $x\in(-1,2.5),\ \epsilon=0.1$.
@@
  • Synthetic training dataset of a target function $f(x)$ constructed by randomly sampling 100 points within a specific range of $x$ from a uniform distribution with some noise level $\epsilon$.
@@
Fig. 3 shows the different single-layer QONN trainings indicating the modes where a photon subtraction is present. The left plot proves that, without subtractions—i.e. QONN with only Gaussian operators—, learning is limited to a linear fit due to the lack of activations. Overall, these results show that adding more neurons—arising from the amount of subtracted modes—in a single layer makes the fitting error tend to zero regardless the complexity of the function, in line with the UAT.
@@ 5.2.1 Moons and circles classification
To test the classification skills of the QONN model, two toy-example tasks to be tackled are the 2D moons and circles classification. In these problems, the target of the network is to learn a nonlinear boundary that tells apart two different classes of points arranged in a continuous two-dimensional grid.
@@
According to the previously defined features, QONNs of $N=2$ modes are sufficient to perform these tasks, since both problems are binary classifications and the input space is two-dimensional and real-valued.
In Fig. 4, the upper row shows how a Gaussian QONN trains up to a linear decision boundary for the moons problem, while a single-layer QONN with one neuron is able to fully discern between the two classes with a curved decision boundary, achieving perfect accuracy. On the other hand, the circles problem needs a radial decision boundary in order to be solved. The lower row of Fig. 4 demonstrates that a single-neuron QONN with two modes fails to capture such a boundary, where a single-neuron QONN with 3 modes instead succeeds with perfect precision. In this last case, adding the third optical mode gives freedom and balance to the output probabilities during the training of the two-category classification since one of the system modes is not observed, that is, the third mode acts as an auxiliary mode that modulates the outcomes of the first two modes whose output values are the ones denoting the probability of belonging to their respective classes.
@@
Figure 4: Moons classification (upper row): two interleaved categories of points. Upper-left: Linear decision boundary with 86.6% accuracy of a Gaussian QONN (no subtractions). Upper-right: S-like decision boundary with 100% of accuracy of a single-layer single-subtraction QONN. Circles classification (lower row): two concentric categories of points. Lower-left: ellipse-like decision boundary with 91.4% of accuracy of a single-layer single-subtraction QONN of 22 modes. Lower-right: radial decision boundary with 100% of accuracy of a single-layer single-subtraction QONN of 33 modes.
@@ 5.2.2 MNIST classification
  • Real-valued inputs: Autoencoder compressing the $28\times 28$ images to a continuous latent space of $5$ real-domain dimensions.
@@
Figure 5: 5-category MNIST classification showing per-class accuracy (opaque lines) and average precision (transparent horizontal lines) for different QONN models with an input pre-processing of 5-dimensional latent space.
@@ 5.3 Non-Gaussian gate synthesis
This admits to train the proposed QONN to match the statistical moments of any complex non-Gaussian gate acting over a particular set of quantum states. This means that the trained QONN block would be able to replicate the action of the chosen non-Gaussian operator upon the target quantum states using physically available quantum optical components, similar to Yukawa et al. (2013).
@@ 6 Discussion
An important conceptual difference between this QONN proposal and earlier continuous-variable neural network architectures Killoran et al. (2019a); Bangar et al. (2023) is the role of Fock-space truncation. In those works, all quantities of interest are ultimately computed in a truncated Fock basis. Formally, this renders the scheme equivalent to a—very large—discrete-variable model: the dynamics is projected onto a finite-dimensional subspace, classical simulation inherits the unfavorable scaling of the truncated Hilbert space, and the resulting figures of merit are not directly expressed in terms of experimentally accessible quadrature statistics. In contrast, the presented approach is based on a genuinely continuous-variable ansatz and an exact evaluation of homodyne moments, without any photon-number cutoff, by combining Bogoliubov transformations with generalized Wick–Isserlis theorems. This provides a closed-form description of the effective activation functions and a classical simulation framework whose cost scales with the number of modes, layers, and photon subtractions, rather than with an artificially imposed Fock cutoff. For the tasks studied here, substantial practical speedup is observed over state-of-the-art Fock-space simulators such as QuTiP and Strawberry Fields.
These differences have direct implications for both expressivity and experimental feasibility. For instance, in the architecture of Bangar et al. (2023), the network depth reaches up to twelve layers with three photon-subtraction events per layer. At realistic squeezing levels and detection efficiencies, such a configuration would lead to heralding rates that are exceedingly low—on the order of one successful detection per day—, implying training times of years even under optimistic assumptions. To keep the numerics tractable, the analysis therefore relies on a photon-number cutoff, typically $n_{\max}=6$, reintroducing the drawbacks of truncation at the simulation level. In the exact framework introduced, by contrast, a single layer with only a few photon-subtracted modes is already sufficient to match or surpass the performance reported Bangar et al. (2023) on comparable tasks, while remaining compatible with realistic heralding probabilities on current photonic platforms. This highlights that a shallow, analytically tractable continuous-variable quantum optical neural network can simultaneously achieve high expressive power, efficient classical simulability, and genuine experimental viability without resorting to Fock space truncation.
@@ 7 Code availability
Since the QONN outputs are computed as sums of traces, and each of these traces can be evaluated independently via the Wick–Isserlis expansion of Eq. (LABEL:perf_match)—which yields a term-wise sum of products—high-performance computing environments are natural candidates for the classical simulation of the model. For this reason, the QuaNNTO library is implemented on top of the high-performance JAX framework Bradbury et al. (2018), together with the Basinhopping method from the SciPy library, which internally uses the L–BFGS–B optimization algorithm with box constraints. In practice, this realizes a hybrid training strategy: an automatic-differentiation (AD) based warm-up stage, implemented purely in JAX, followed by a quasi-Newton optimization method (L–BFGS–B) acting as a local refiner. This combination enables large-scale simulations of gradient-based QONN training procedures on distributed and parallel classical hardware.
@@ Appendix A Nonlinearity of the creation operator in phase-space
Eq. (A2) is equivalent to Laguerre polynomials defined as creation operators over coherent states [1], whose expansion leads to the general expression
@@
$\langle\hat{x}(\alpha,r)\rangle=\sqrt{2}\Re\Big(\underbrace{e^{r}\,\alpha}_{\text{Linear map of }\alpha}\;+\;\underbrace{\frac{\alpha\cdot e^{2r}\cosh r}{\alpha^{2}\cdot e^{2r}+\cosh^{2}r}}_{\text{Rational bump of }\alpha}\Big).$ (A6)
Figure 7: Nonlinearity induced by creation operator in $\hat{x}$ quadrature of squeezed-coherent states with respect to their real displacements $\alpha\in(-2,2)$. Left: $\langle\hat{x}(\alpha)\rangle_{n}$ from Eq. (A3) for $n=0..4$ photon additions. Center and right: $\langle\hat{x}(\alpha,r)\rangle$ from Eq. (A6) for different values of squeezing in momentum (center, $r\geq 0$) and position (right, $r\leq 0$).
From the explicit shapes of Eq. (A3) shown in the left plot of Fig. 7, one can clearly notice the smooth nonlinear sigmoid-like effect of the photon addition operations—bounded and non-polynomial—over a real coherent state when the displacement is relatively small, even without squeezing. However, the effect tends to be linear when the displacement is more pronounced, making it a saturated nonlinearity.
The effect of the photon addition is similar when the initial coherent state is squeezed previously to the addition, as seen in Eq. (A6). The activation function is also non-polynomial and smooth. The center plot of Fig. 7 shows how squeezing momentum quadrature pronounces the nonlinearity induced by the photon creation in the position quadrature mean, while squeezing position (right plot) tends to linearize such effect when the—absolute value—squeezing factor enlarges.
@@ Appendix B Exact classical simulation of the CV-QONN model
To classically simulate the QONN from Fig. 1 using the Wick–Isserlis expansion of Eq. (LABEL:perf_match), all the Gaussian operators from the layers have to be dragged to the beginning of the circuit for the state to remain Gaussian and, afterwards, apply all the non-Gaussianity which would become a combination of creation and annihilation operators.
This is achieved by the Bogoliubov unitary action of the general Gaussian operator $\hat{G}$ of the form of Eq. (9) over the annihilation operator [8]
@@
$\rho_{NG}=\frac{\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\ \rho_{G}\ \hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}}{\text{Tr}\big[\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\ \rho_{G}\ \hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}\big]},$ (B7)
@@
$\begin{split}\tilde{f}(\bm{\alpha})=\text{Tr}\left[\hat{x}_{1}\rho_{NG}\right]=\frac{\text{Tr}\left[\frac{(\hat{a}_{1}+\hat{a}^{\dagger}_{1})}{\sqrt{2}}\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\ \rho_{G}\ \hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}\right]}{\text{Tr}\big[\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\ \rho_{G}\ \hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}\big]}=\frac{\sqrt{2}\cdot\Re\Big(\text{Tr}\left[\hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}\hat{a}_{1}\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\rho_{G}\right]\Big)}{\text{Tr}\big[\hat{\mathbb{S}}_{1}^{(2)\dagger}\hat{a}_{1}^{\dagger}\hat{a}_{1}\hat{\mathbb{S}}_{1}^{(2)}\rho_{G}\big]},\end{split}$ (B9)
@@
Each of these trace expressions has the form of Eq. (LABEL:perf_match) and is computed by Wick–Isserlis using the set $\mathcal{P}$ of all loop perfect matchings of the ladder operators in the expression. For an expression with $M$ ladder operators, the total number of different matchings increases combinatorially guided by Eq. (11).
Note that, additionally, this needs to be applied for the normalization factor of the QONN’s final quantum state $\rho_{NG}$ of Eq. (B7) with $O=1$ in Eq. (B10).